OpenStudy (celinegirl):

check my answer - A Jack invested some money in a bank at a fixed rate of interest compounded annually. The equation below shows the value of his investment after x years: f(x) = 300(1.02)^x What was the average rate of change of the value of Jack's investment from the third year to the fifth year? 6.43 dollars per year 8.24 dollars per year 12.86 dollars per year 14.26 dollars per year

OpenStudy (welshfella):

Find the value of the investment after 3 years and also the value after 5 years. Do you know how to do that?

OpenStudy (mathmale):

the proper way to write your "f(x) = 300(1.02)x" is \[f(x) = 300(1.02)^x\] that x is an exponent, not a multiplier. average rate of change:\[\frac{ f(5)-f(3) }{ 5-3 }\] Please evaluate the function f(x) at time 3 and time 5.

OpenStudy (celinegirl):

forgot to add that. oops.. I'm just asking if my answer is correct

OpenStudy (kobeni-chan):

@celinegirl I can confirm if your answer is correct or not if you post the process you took to get that answer. Otherwise, it may seem like someone's just asking if their guess is right, you know what I mean?

OpenStudy (celinegirl):

Yes but I did not guess.. i do most of my work on paper

OpenStudy (kobeni-chan):

Ok, can you type it up for us?